# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ III.6 Choice of source functions $Z_{R}$, $Z_{S}$
$\left(\widetilde{T}_{vv}\right)_{,u}=e^{S}L_{0}\left[\left(\varepsilon_{0}M^{-5}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,v}-\frac{2Q}{R^{\frac{3}{2}}}Q_{,v}-\varepsilon_{0}M^{-3}\frac{r_{,v}}{M}Z_{Rr}\right]$ (III.34)
From Eq. (III.33) we can see that if all sources are zero and in addition $Q_{,u}=0$, then $\widetilde{T}_{uu,v}=0$. Since $T_{uu}$ vanishes on the initial hypersurface, it follows that in such a case, $T_{uu}$ vanishes in the entire domain of dependence. (Also notice that the parameter $Z_{R0}$ does not affect the conservation of either $\widetilde{T}_{uu}$ or $\widetilde{T}_{vv}$.)
As was mentioned in the Introduction, the analytical approximation is based on a key assumption about the behavior of the RSET in the evolving geometry. Here we shall specify this assumption. In the near-IH RN background (in Unruh state) it was found that both $Z_{R}$ and $Z_{S}$ [32] vanish sufficiently fast in both $u$ and $v$ and hence become negligibly small on approaching the IH (see also footnote II). Our key assumption is that in the evolving geometry too, these source functions remain negligibly small. This assumption leads to conservation of the fluxes $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$ in the near-IH region: $\widetilde{T}_{uu,v}=0$ and $\widetilde{T}_{vv,u}=0$. To see this, we evaluate $\widetilde{T}_{uu,v}$ by taking the $v$ derivative of Eq. (II.44) and also substituting the evolution equations Eqs. (II.117,II.118):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (III.35)
$=\left(e^{S}L_{0}F_{1}(R)+Z_{R}\right)S_{,u}$ (III.36)
$+\left(e^{S}L_{0}F_{2}(R)+Z_{S}\right)R_{,u}$ (III.37)
$-\left(e^{S}L_{0}F_{1}(R)+Z_{R}\right)_{,u}$ (III.38)
Since $e^{S}$ decays exponentially with $u$ and $v$ (see subsection IV.5), the quick vanishing of $Z_{R}$ and $Z_{S}$ leads to $\left(\widetilde{T}_{uu}\right)_{,v}=0$ in the near-IH region. The same argument also applies to $\left(\widetilde{T}_{vv}\right)_{,u}$.
IV Analytical approximation for the near-IH metric inside slowly evaporating black holes
@@
$\varepsilon=M_{,v}$ (IV.1)
is small, $\varepsilon\ll 1$. We can build an approximation assuming $\varepsilon$ is small.
IV.1 Regions inside black hole
Figure IV.1: metrics inside black hole
We divide the space-time domain inside the black hole to three regions, as seen in Figure IV.1 . Those regions, numbered 1-3, are ordered in chronological order. Each two successive regions overlap in their area of validity, and together they cover the entire interior of the black hole up to the inner horizon.
Region 1 is the neighborhood of the event horizon. This is taken here to be ingoing charged Vaidya metric Eq. (II.100), which we later convert into double-null coordinates.
@@ IV.1 Regions inside black hole
IV.2 Basic ansatz of the analytical approximation for region 3
Consider first the near-IH domain in region 2 where the geometry is locally RN. We saw in Eq. (II.90) that in RN $S_{,u}=S_{,v}=-\kappa_{-}$ near the IH. We therefore get that $S$ is monotonically decreasing in region 2, and becomes very negative upon increasing $u$ and $v$. Therefore, $e^{S}$ becomes negligibly small at large $u+v$.
@@ IV.2 Basic ansatz of the analytical approximation for region 3
Using the evolution equations Eq. (II.117,II.118) and neglecting the factor $e^{S}$ – and also assuming negligibly small sources $Z_{R}\approx Z_{S}\approx 0$ (see subsection III.6) – we get, throughout region 3:
$R_{,uv}=0,\quad S_{,uv}=0$ (IV.2)
@@
Figure IV.2: Diagram explaining analytical approximation
@@
IV.3 Computing $S_{,v}$
Refer to Figure IV.2. Consider a point $P$ deep in region 3. We first focus on computing $S_{,v}$ at point $P$. Since $S_{,v}(v)$ is constant along the red constant-$v$ ray, we can read its value from point $P_{v}$, which is in the overlap of regions 2 and 3, and therefore its value is obtained from the corresponding expression in region 2 near IH (see Eq. (II.90)):
$S_{,v}=-\kappa_{-}^{v}(v)$ (IV.5)
@@ IV.3 Computing $S_{,v}$
$\kappa_{-}^{v}(v)\equiv\kappa_{-}(M(v),Q^{v}(v))$ (IV.6)
IV.4 Computing $S_{,u}$
@@
$S_{,u}(P)=-\kappa_{-}(P_{u})$ (IV.7)
Here $\kappa_{-}(P_{u})$ denotes the value of $\kappa_{-}$ at the local RN at point $P_{u}$. We specifically pick the point $P_{u}$ such that it is located at $u+v$ of order, say, $10^{2}M$. At that point, $e^{S}$ is already negligible, and yet $\varepsilon\cdot\left(v+u\right)$ is also negligible. (These two properties are valid throughout the overlap of region 2 and 3.) As mentioned above, the local RN at a given point in region 2 is the same along the constant-$v$ ray that passes through that point. We could therefore mark a constant-$v$ ray from this point $P_{u}$ up to the point $P_{2}$ on the EH. However we prefer to pick another point $P_{1}$ on the EH to represent the local RN at point $P_{u}$. This point $P_{1}$ is obtained by continuing from point $P_{u}$ along the dashed blue constant-$u$ ray until we meet the hypersurface $r=M$ (which is also the $r_{*}=0$ line) at point $P_{0}$. Then we move from $P_{0}$ along the dashed purple constant-$v$ ray to reach the point $P_{1}$ on the EH. We prefer to work with the representative point $P_{1}$ rather than $P_{2}$ because this choice leads to a significantly simpler expression for $S_{,u}$ (and $R_{,u}$) in region 3. However, it remains to be explained why this choice, namely the replacement of point $P_{2}$ by point $P_{1}$, is justified.
@@
$\kappa_{-}^{u}(u)\equiv\kappa_{-}(M^{u}(u),Q^{u}(u))$ (IV.11)
@@
$\kappa_{-}^{u}(u)\equiv\kappa_{-}(v=-u)$ (IV.12)
@@
$\kappa_{+}^{u}(u)\equiv\kappa_{+}(v=-u)$ (IV.13)
@@
$\kappa_{-}(P_{u})=\kappa_{-}^{u}(u_{P_{u}})=\kappa_{-}^{u}(u_{P})$ (IV.14)
We can therefore rewrite Eq. (IV.7) as
$S_{,u}(P)=-\kappa_{-}^{u}(u_{P})$ (IV.15)
@@
$S_{,u}=-\kappa_{-}^{u}(u)$ (IV.16)
IV.5 Justifying the basic ansatz
Recalling Eqs. (IV.5,IV.16) and the fact that $\kappa_{-}>0$, we see that in region 3, $S$ continues to decrease monotonically with both $u$ and $v$. Since $e^{S}$ was already exponentially small in the overlap between region 2 and region 3, it is totally negligible throughout region 3 – confirming our former ansatz about this issue.
Correspondingly we can neglect the factor $e^{S}$ also in (III.33,III.34) and we get that:
$\widetilde{T}_{uu,v}=\widetilde{T}_{vv,u}=0$ (IV.17)
@@ IV.5 Justifying the basic ansatz
IV.6 Computing $R_{,v}$
We next focus on $R_{,v}$ in region 3. We recall that $\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}=R_{,v}S_{,v}-R_{,vv}$, and we make an ansatz that $R_{,vv}\ll R_{,v}S_{,v}$ (to be justified later) and therefore, using Eq. (IV.5) and neglecting $R_{,vv}$, we get:
$R_{,v}=\frac{\widetilde{T}_{vv}}{S_{,v}}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}$ (IV.18)
@@ IV.6 Computing $R_{,v}$
$F=M^{m}\varepsilon^{n}H_{F}\left(\frac{Q^{v}(v)}{M(v)}\right)$ (IV.19)
@@
$F_{,v}=M^{m-1}\varepsilon^{n+1}H_{F^{\prime}}\left(\frac{Q^{v}(v)}{M(v)}\right)$ (IV.20)
@@
$\left(\frac{Q^{v}}{M}\right)_{,v}=-\frac{Q^{v}M_{,v}}{M^{2}}+\frac{Q_{,v}^{v}}{M}$ (IV.21)
From semiclassical theory we know that Eq. (III.7) (setting $M^{(+)}\rightarrow M$ and $Q\rightarrow Q^{v}$):
$M_{,v}\equiv-\varepsilon=-\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q^{v}}{M}\right)$ (IV.22)
@@
$Q_{,v}^{v}=-\frac{\hbar}{M^{2}}H_{Q}\left(\frac{Q^{v}}{M}\right)=-\varepsilon\frac{H_{Q}\left(\frac{Q^{v}}{M}\right)}{H_{M}\left(\frac{Q^{v}}{M}\right)}\equiv-\varepsilon H_{q}\left(\frac{Q^{v}}{M}\right)$ (IV.23)
@@
$\left(\frac{Q^{v}}{M}\right)_{,v}=\frac{\varepsilon Q^{v}}{M^{2}}-\frac{\varepsilon H_{q}\left(\frac{Q^{v}}{M}\right)}{M}\equiv\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (IV.24)
@@
$\frac{d}{dv}H\left(\frac{Q^{v}}{M}\right)=H^{\prime}\left(\frac{Q^{v}}{M}\right)\frac{d}{dv}\left(\frac{Q^{v}}{M}\right)=H^{\prime}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (IV.25)
where $H^{\prime}$ denotes the derivative of the function $H$ with respect to its argument $\frac{Q^{v}}{M}$. We next calculate $\varepsilon_{,v}$ by differentiating Eq. (IV.22):
$\varepsilon_{,v}$ $=-2\frac{\hbar M_{,v}}{M^{3}}H_{M}\left(\frac{Q^{v}}{M}\right)+\frac{\hbar}{M^{2}}\frac{d}{dv}H_{M}\left(\frac{Q^{v}}{M}\right)$ (IV.26)
$=-2\frac{\hbar M_{,v}}{M^{3}}H_{M}\left(\frac{Q^{v}}{M}\right)+\frac{\hbar}{M^{2}}H^{\prime}_{M}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (IV.27)
$=\frac{\hbar}{M^{2}}\frac{\varepsilon}{M}\left[2H_{M}\left(\frac{Q^{v}}{M}\right)+H^{\prime}_{M}\left(\frac{Q^{v}}{M}\right)H_{v}\left(\frac{Q^{v}}{M}\right)\right]$ (IV.28)
$\equiv\frac{\varepsilon^{2}}{M}H_{\varepsilon}\left(\frac{Q^{v}}{M}\right)$ (IV.29)
where $H^{\prime}_{M}$ denotes the derivative of the function $H_{M}$ with respect to its argument $\frac{Q^{v}}{M}$. Finally, we take the derivative of $F$ in Eq. (IV.19) to get:
$F_{,v}$ $=mM_{,v}M^{m-1}\varepsilon^{n}H_{F}\left(\frac{Q^{v}}{M}\right)+n\varepsilon_{,v}M^{m}\varepsilon^{n-1}H_{F}\left(\frac{Q^{v}}{M}\right)+M^{m}\varepsilon^{n}H^{\prime}_{F}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (IV.30)
$=-M^{m-1}\varepsilon^{n+1}mH_{F}\left(\frac{Q^{v}}{M}\right)+M^{m-1}\varepsilon^{n+1}nH_{\varepsilon}\left(\frac{Q^{v}}{M}\right)H_{F}\left(\frac{Q^{v}}{M}\right)$ (IV.31)
$+M^{m-1}\varepsilon^{n+1}H^{\prime}_{F}\left(\frac{Q^{v}}{M}\right)H_{v}\left(\frac{Q^{v}}{M}\right)$ (IV.32)
$\equiv M^{m-1}\varepsilon^{n+1}H_{F^{\prime}}\left(\frac{Q^{v}}{M}\right)$ (IV.33)
We proved Eq. (IV.20). Now we will use this result for Eq. (IV.18). First, recall from Eq. (III.17) that $\widetilde{T}_{vv}^{(-)}$ can be expressed as:
$\widetilde{T}_{vv}^{(-)}=\varepsilon H_{T_{vv}}^{(-)}\left(\frac{Q^{v}}{M}\right)$ (IV.34)
Secondly, recall that for $\kappa_{-},$Eq. (II.54) can be written as:
$\kappa_{-}=\frac{1}{M}H_{\kappa}\left(\frac{Q^{v}}{M}\right)$ (IV.35)
@@
$R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}=-\varepsilon M\frac{H_{T_{vv}}^{(-)}\left(\frac{Q^{v}}{M}\right)}{H_{\kappa}\left(\frac{Q^{v}}{M}\right)}\equiv\varepsilon MH_{R}\left(\frac{Q^{v}}{M}\right)$ (IV.36)
We can now apply Eq. (IV.20) to $F=R_{,v}$ to get $R_{,vv}$:
$R_{,vv}=\varepsilon^{2}H_{vv}\left(\frac{Q^{v}}{M}\right)$ (IV.37)
@@
$R_{,v}S_{,v}$ $=-R_{,v}\kappa_{-}=\left[\varepsilon MH_{R}\left(\frac{Q^{v}}{M}\right)\right]\left[\frac{1}{M}H_{\kappa}\left(\frac{Q^{v}}{M}\right)\right]$ (IV.38)
$=\varepsilon H_{RS}\left(\frac{Q^{v}}{M}\right)$ (IV.39)

### v2

@@ 3.6 Choice of source functions $Z_{R}$, $Z_{S}$
$\left(\widetilde{T}_{vv}\right)_{,u}=e^{S}L_{0}\left[\left(\varepsilon_{0}M^{-5}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,v}-\frac{2Q}{R^{\frac{3}{2}}}Q_{,v}-\varepsilon_{0}M^{-3}\frac{r_{,v}}{M}Z_{Rr}\right]$ (3.34)
From Eq. (3.33) we can see that if all sources are zero and in addition $Q_{,u}=0$, then $\widetilde{T}_{uu,v}=0$. Since $T_{uu}$ vanishes on the initial hypersurface, it follows that in such a case, $T_{uu}$ vanishes in the entire domain of dependence. (Also notice that the parameter $Z_{R0}$ does not affect the conservation of either $\widetilde{T}_{uu}$ or $\widetilde{T}_{vv}$.)
As was mentioned in the Introduction, the analytical approximation is based on a key assumption about the behavior of the RSET in the evolving geometry. Here we shall specify this assumption. In the near-IH RN background (in Unruh state) it was found that both $Z_{R}$ and $Z_{S}$ [30] vanish sufficiently fast in both $u$ and $v$ and hence become negligibly small on approaching the IH (see also footnote 5). Our key assumption is that in the evolving geometry too, these source functions remain negligibly small. This assumption leads to conservation of the fluxes $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$ in the near-IH region: $\widetilde{T}_{uu,v}=0$ and $\widetilde{T}_{vv,u}=0$. To see this, we evaluate $\widetilde{T}_{uu,v}$ by taking the $v$ derivative of Eq. (2.43) and also substituting the evolution equations Eqs. (2.115,2.116):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (3.35)
$=\left(e^{S}L_{0}F_{1}(R)+Z_{R}\right)S_{,u}$ (3.36)
$+\left(e^{S}L_{0}F_{2}(R)+Z_{S}\right)R_{,u}$ (3.37)
$-\left(e^{S}L_{0}F_{1}(R)+Z_{R}\right)_{,u}$ (3.38)
Since $e^{S}$ decays exponentially with $u$ and $v$ (see subsection 4.5), the quick vanishing of $Z_{R}$ and $Z_{S}$ leads to $\left(\widetilde{T}_{uu}\right)_{,v}=0$ in the near-IH region. The same argument also applies to $\left(\widetilde{T}_{vv}\right)_{,u}$.
4 Analytical approximation for the near-IH metric inside slowly evaporating black holes
@@
$\varepsilon=M_{,v}$ (4.1)
is small, $\varepsilon\ll 1$. We can build an approximation assuming $\varepsilon$ is small. Note that this $\varepsilon$ is distinct from $\varepsilon_{0}$: $\varepsilon$ depends on $v$ while $\varepsilon_{0}$ is a constant associated with the substitution $v=v_{0}$. This difference reflects the fact that at very late $v$, $\varepsilon$ is no longer small, being proportional to $M(v)^{-2}$.
4.1 Regions inside black hole
Figure 4.1: metrics inside black hole
We divide the space-time domain inside the black hole to three regions, as seen in Figure 4.1 . Those regions, numbered 1-3, are ordered in chronological order. Each two successive regions overlap in their area of validity, and together they cover the entire interior of the black hole up to the inner horizon.
Region 1 is the neighborhood of the event horizon. This is taken here to be ingoing charged Vaidya metric Eq. (2.99), which we later convert into double-null coordinates.
@@ 4.1 Regions inside black hole
4.2 Basic ansatz of the analytical approximation for region 3
Consider first the near-IH domain in region 2 where the geometry is locally RN. We saw in Eq. (2.89) that in RN $S_{,u}=S_{,v}=-\kappa_{-}$ near the IH. We therefore get that $S$ is monotonically decreasing in region 2, and becomes very negative upon increasing $u$ and $v$. Therefore, $e^{S}$ becomes negligibly small at large $u+v$.
@@ 4.2 Basic ansatz of the analytical approximation for region 3
Using the evolution equations Eq. (2.115,2.116) and neglecting the factor $e^{S}$ – and also assuming negligibly small sources $Z_{R}\approx Z_{S}\approx 0$ (see subsection 3.6) – we get, throughout region 3:
$R_{,uv}=0,\quad S_{,uv}=0$ (4.2)
@@
Figure 4.2: Diagram explaining analytical approximation
@@
4.3 Computing $S_{,v}$
Refer to Figure 4.2. Consider a point $P$ deep in region 3. We first focus on computing $S_{,v}$ at point $P$. Since $S_{,v}(v)$ is constant along the red constant-$v$ ray, we can read its value from point $P_{v}$, which is in the overlap of regions 2 and 3, and therefore its value is obtained from the corresponding expression in region 2 near IH (see Eq. (2.89)):
$S_{,v}=-\kappa_{-}^{v}(v)$ (4.5)
@@ 4.3 Computing $S_{,v}$
$\kappa_{-}^{v}(v)\equiv\kappa_{-}(M(v),Q^{v}(v))$ (4.6)
4.4 Computing $S_{,u}$
@@
$S_{,u}(P)=-\kappa_{-}(P_{u})$ (4.7)
Here $\kappa_{-}(P_{u})$ denotes the value of $\kappa_{-}$ at the local RN at point $P_{u}$. We pick the point $P_{u}$ in the matching zone between region 2 and region 3. In this matching zone, both approximation of region 2 and region 3 should apply. The approximation of region 3 requires $e^{S}$ to be negligible. The region 2 approximation requires that the near IH drift in $R$ is still negligible, that is $\varepsilon\cdot\left(v+u\right)\ll M$. We choose this matching zone to be at the region where $u+v$ is of order, say, $\frac{1}{\kappa_{-}}\log\varepsilon$. At that point, $e^{S}$ is indeed negligible, because $S\approx-\kappa_{-}(u+v)$ so $e^{S}\approx\varepsilon\ll 1$. In addition, $\varepsilon\cdot\left(v+u\right)=-\frac{1}{\kappa_{-}}\varepsilon\log\varepsilon$ is negligible compared to $M$ in that zone.
As mentioned above, the local RN at a given point in region 2 is the same along the constant-$v$ ray that passes through that point. We could therefore mark a constant-$v$ ray from this point $P_{u}$ up to the point $P_{2}$ on the EH. However we prefer to pick another point $P_{1}$ on the EH to represent the local RN at point $P_{u}$. This point $P_{1}$ is obtained by continuing from point $P_{u}$ along the dashed blue constant-$u$ ray until we meet the hypersurface $r=M$ (which is also the $r_{*}=0$ line) at point $P_{0}$. Then we move from $P_{0}$ along the dashed purple constant-$v$ ray to reach the point $P_{1}$ on the EH. We prefer to work with the representative point $P_{1}$ rather than $P_{2}$ because this choice leads to a significantly simpler expression for $S_{,u}$ (and $R_{,u}$) in region 3. However, it remains to be explained why this choice, namely the replacement of point $P_{2}$ by point $P_{1}$, is justified.
@@
$\kappa_{-}^{u}(u)\equiv\kappa_{-}(M^{u}(u),Q^{u}(u))$ (4.11)
@@
$\kappa_{-}^{u}(u)\equiv\kappa_{-}(v=-u)$ (4.12)
@@
$\kappa_{+}^{u}(u)\equiv\kappa_{+}(v=-u)$ (4.13)
@@
$\kappa_{-}(P_{u})=\kappa_{-}^{u}(u_{P_{u}})=\kappa_{-}^{u}(u_{P})$ (4.14)
We can therefore rewrite Eq. (4.7) as
$S_{,u}(P)=-\kappa_{-}^{u}(u_{P})$ (4.15)
@@
$S_{,u}=-\kappa_{-}^{u}(u)$ (4.16)
4.5 Justifying the basic ansatz
Recalling Eqs. (4.5,4.16) and the fact that $\kappa_{-}>0$, we see that in region 3, $S$ continues to decrease monotonically with both $u$ and $v$. Since $e^{S}$ was already exponentially small in the overlap between region 2 and region 3, it is totally negligible throughout region 3 – confirming our former ansatz about this issue.
Correspondingly we can neglect the factor $e^{S}$ also in (3.33,3.34) and we get that:
$\widetilde{T}_{uu,v}=\widetilde{T}_{vv,u}=0$ (4.17)
@@ 4.5 Justifying the basic ansatz
4.6 Computing $R_{,v}$
We next focus on $R_{,v}$ in region 3. We recall that $\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}=R_{,v}S_{,v}-R_{,vv}$, and we make an ansatz that $R_{,vv}\ll R_{,v}S_{,v}$ (to be justified later) and therefore, using Eq. (4.5) and neglecting $R_{,vv}$, we get:
$R_{,v}=\frac{\widetilde{T}_{vv}}{S_{,v}}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}$ (4.18)
@@ 4.6 Computing $R_{,v}$
$F=M^{m}\varepsilon^{n}H_{F}\left(\frac{Q^{v}(v)}{M(v)}\right)$ (4.19)
@@
$F_{,v}=M^{m-1}\varepsilon^{n+1}H_{F^{\prime}}\left(\frac{Q^{v}(v)}{M(v)}\right)$ (4.20)
@@
$\left(\frac{Q^{v}}{M}\right)_{,v}=-\frac{Q^{v}M_{,v}}{M^{2}}+\frac{Q_{,v}^{v}}{M}$ (4.21)
From semiclassical theory we know that Eq. (3.7) (setting $M^{(+)}\rightarrow M$ and $Q\rightarrow Q^{v}$):
$M_{,v}\equiv-\varepsilon=-\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q^{v}}{M}\right)$ (4.22)
@@
$Q_{,v}^{v}=-\frac{\hbar}{M^{2}}H_{Q}\left(\frac{Q^{v}}{M}\right)=-\varepsilon\frac{H_{Q}\left(\frac{Q^{v}}{M}\right)}{H_{M}\left(\frac{Q^{v}}{M}\right)}\equiv-\varepsilon H_{q}\left(\frac{Q^{v}}{M}\right)$ (4.23)
@@
$\left(\frac{Q^{v}}{M}\right)_{,v}=\frac{\varepsilon Q^{v}}{M^{2}}-\frac{\varepsilon H_{q}\left(\frac{Q^{v}}{M}\right)}{M}\equiv\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.24)
@@
$\frac{d}{dv}H\left(\frac{Q^{v}}{M}\right)=H^{\prime}\left(\frac{Q^{v}}{M}\right)\frac{d}{dv}\left(\frac{Q^{v}}{M}\right)=H^{\prime}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.25)
where $H^{\prime}$ denotes the derivative of the function $H$ with respect to its argument $\frac{Q^{v}}{M}$. We next calculate $\varepsilon_{,v}$ by differentiating Eq. (4.22):
$\varepsilon_{,v}$ $=-2\frac{\hbar M_{,v}}{M^{3}}H_{M}\left(\frac{Q^{v}}{M}\right)+\frac{\hbar}{M^{2}}\frac{d}{dv}H_{M}\left(\frac{Q^{v}}{M}\right)$ (4.26)
$=-2\frac{\hbar M_{,v}}{M^{3}}H_{M}\left(\frac{Q^{v}}{M}\right)+\frac{\hbar}{M^{2}}H^{\prime}_{M}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.27)
$=\frac{\hbar}{M^{2}}\frac{\varepsilon}{M}\left[2H_{M}\left(\frac{Q^{v}}{M}\right)+H^{\prime}_{M}\left(\frac{Q^{v}}{M}\right)H_{v}\left(\frac{Q^{v}}{M}\right)\right]$ (4.28)
$\equiv\frac{\varepsilon^{2}}{M}H_{\varepsilon}\left(\frac{Q^{v}}{M}\right)$ (4.29)
where $H^{\prime}_{M}$ denotes the derivative of the function $H_{M}$ with respect to its argument $\frac{Q^{v}}{M}$. Finally, we take the derivative of $F$ in Eq. (4.19) to get:
$F_{,v}$ $=mM_{,v}M^{m-1}\varepsilon^{n}H_{F}\left(\frac{Q^{v}}{M}\right)+n\varepsilon_{,v}M^{m}\varepsilon^{n-1}H_{F}\left(\frac{Q^{v}}{M}\right)+M^{m}\varepsilon^{n}H^{\prime}_{F}\left(\frac{Q^{v}}{M}\right)\frac{\varepsilon}{M}H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.30)
$=-M^{m-1}\varepsilon^{n+1}mH_{F}\left(\frac{Q^{v}}{M}\right)+M^{m-1}\varepsilon^{n+1}nH_{\varepsilon}\left(\frac{Q^{v}}{M}\right)H_{F}\left(\frac{Q^{v}}{M}\right)$ (4.31)
$+M^{m-1}\varepsilon^{n+1}H^{\prime}_{F}\left(\frac{Q^{v}}{M}\right)H_{v}\left(\frac{Q^{v}}{M}\right)$ (4.32)
$\equiv M^{m-1}\varepsilon^{n+1}H_{F^{\prime}}\left(\frac{Q^{v}}{M}\right)$ (4.33)
We proved Eq. (4.20). Now we will use this result for Eq. (4.18). First, recall from Eq. (3.17) that $\widetilde{T}_{vv}^{(-)}$ can be expressed as:
$\widetilde{T}_{vv}^{(-)}=\varepsilon H_{T_{vv}}^{(-)}\left(\frac{Q^{v}}{M}\right)$ (4.34)
Secondly, recall that for $\kappa_{-},$Eq. (2.53) can be written as:
$\kappa_{-}=\frac{1}{M}H_{\kappa}\left(\frac{Q^{v}}{M}\right)$ (4.35)
@@
$R_{,v}=-\frac{\widetilde{T}_{vv}^{(-)}}{\kappa_{-}(v)}=-\varepsilon M\frac{H_{T_{vv}}^{(-)}\left(\frac{Q^{v}}{M}\right)}{H_{\kappa}\left(\frac{Q^{v}}{M}\right)}\equiv\varepsilon MH_{R}\left(\frac{Q^{v}}{M}\right)$ (4.36)
We can now apply Eq. (4.20) to $F=R_{,v}$ to get $R_{,vv}$:
$R_{,vv}=\varepsilon^{2}H_{vv}\left(\frac{Q^{v}}{M}\right)$ (4.37)
@@
$R_{,v}S_{,v}$ $=-R_{,v}\kappa_{-}=\left[\varepsilon MH_{R}\left(\frac{Q^{v}}{M}\right)\right]\left[\frac{1}{M}H_{\kappa}\left(\frac{Q^{v}}{M}\right)\right]$ (4.38)
$=\varepsilon H_{RS}\left(\frac{Q^{v}}{M}\right)$ (4.39)